40. In testing H0: μ1 − μ2 = 5 vs. Ha: μ1 − μ2 > 5, the test statistic value z is found to be 1.69. What is the p-value of the test?

A: 0.0910

B: 0.0455

C: 0.3023

D: 0.1977

41. When testing H0: μ1 − μ2 = 0 vs. H1: μ1 − μ2 < 0, the observed value of the z-score was found to be −2.15. What would the p-value for this test be?

A: 0.0316

B: 0.0158

C: 0.9842

D: 0.9684
Answer #1

solution:

right tail

z= 1.69

P(z >1.69 ) = 1 - P(z < 1.69 ) = 1-0.9545=0.0455 using z table

P-value = 0.0455

b.

left tail

z=-2.15

P(z < -2.15 ) = 0.0158 using z table

P-value = 0.0158